OpenStudy (anonymous):

lecture 1:rate of change calculating the area prof states that 2XoYo=2 not sure how he got this can you help? thanks in advance

OpenStudy (anonymous):

Because Yo=(1/Xo), 2XoYo=2Xo(1/Xo) which cancels to equal 2. You can see this worked out on pg 4 of the lecture notes for lecture 1.

OpenStudy (anonymous):

And Yo=(1/Xo) because the equation for that example was y=1/x